OpenStudy (anonymous):

Could someone help me on this question, Find the derivative of g(x) = 2x/5x-1

OpenStudy (jamesj):

Let u = 2x and v = 5x - 1. Then use the quotient rule on your function g = u/v

OpenStudy (jamesj):

Do you know the quotient rule for differentiation?

OpenStudy (anonymous):

um somewhat, my prof hasn't really taught us it yet but gave us this question to try it out

OpenStudy (jamesj):

\[g' = \frac{v.u' - u.v'}{v^2}\] So here u' = 2 v' = 5 So now substitute into the quotient rule and simplify.

OpenStudy (anonymous):

Ohh okay, thanks a lot i appreciate it.